OpenStudy (anonymous):

What is a possible value for the missing term of the geometric sequence? 50, ___, 450,

OpenStudy (tkhunny):

Really Big Hint: 450 / 50 = 9

OpenStudy (anonymous):

would the answer be 150?

OpenStudy (tkhunny):

Why would that be a good answer? Is there a common factor? 150/50 = 3 450/150 = 3 I think you have it!

OpenStudy (anonymous):

yay! thank you so much :)